### v2

Abstract
In the present work, we extend the $f(Q)$ symmetric teleparallel gravity by introducing an arbitrary coupling between the non-metricity $Q$ and matter Lagrangian $L_{m}$ in the Lagrangian density $f$ of the theory, which thus leads to the $f\left(Q,L_{m}\right)$ theory. This generalisation encompasses Coincident General Relativity (CGR), and the Symmetric Teleparallel Equivalent to GR (STEGR). Using the metric formalism, we derive the field equation of the theory, which generalizes the field equations of $f(Q)$ gravity. From the study of the covariant divergence of the field equations, it follows that the presence of the geometry-matter coupling leads to the non-conservation of the matter energy-momentum tensor. The cosmological implications of the theory are investigated in the case of a flat, homogeneous, and isotropic Friedmann-Lemaitre-Robertson-Walker geometry. As a first step in this direction, we obtain the modified Friedmann equations for the $f(Q,L_{m})$ gravity in a general form. Specific cosmological models are investigated for several choices of $f(Q,L_{m})$, including $f(Q,L_{m})=-\alpha Q+2L_{m}+\beta$, and $f(Q,L_{m})=-\alpha Q+(2L_{m})^{2}+\beta$, respectively. Comparative analyses with the standard $\Lambda$ CDM paradigm are carried out, and the observational implications of the models are investigated in detail.
1 Introduction
The theory of General Relativity (GR) [1, 2, 3] revolutionized our understanding of gravity by conceptualizing it not as a conventional force but as an inherent property of spacetime, rooted in Riemannian geometry [4]. Thus, General Relativity developed Contrary to Einstein’s initial intent of not “geometrizing gravity” [5]. In GR, the metric and matter interact minimally, as defined by the Einstein-Hilbert action ($S_{EH}$), given as $S_{EH}=(1/2\kappa)\int\sqrt{-g}R\,d^{4}x+S_{m}$, where $\kappa$ denotes the gravitational coupling constant term. $g$ is the determinant of the metric tensor, and $R$ represents the Ricci scalar. By $S_{m}$, we have denoted the matter action. This linearity in $R$ gives a second-order field equation $G_{\mu\nu}=R_{\mu\nu}-(1/2)R\,g_{\mu\nu}=\kappa T_{\mu\nu}$, which governs the dynamics of matter in the curved spacetime and relates geometry, described by the Einstein tensor $G_{\mu\nu}$ to the matter energy-momentum tensor $T_{\mu\nu}$.
A large number of observations have established GR as a very successful theory of gravity by confirming many of its predictions, such as the deflection of light by the Sun’s gravitational field [6], the perihelion motion of Mercury [7], the existence of gravitational waves [8], gravitational redshift [9], orbital decay of the Hulse-Taylor binary pulsar [10], and the radar echo delay [11, 12], respectively. An in-depth analysis of all the experimental and observational tests of GR can be found in [13, 14].
Despite its remarkable success, which spanned almost one hundred years, the theory of General Relativity currently faces numerous challenges. At a quantum level, it cannot explain the quantum properties of the gravitational interaction. Also, gravitational collapse can result in geodesic incompleteness under specific assumptions regarding the energy-momentum tensor. This implies that certain types of geodesics are constrained by an upper limit on an affine parameter, indicating a singular structure in spacetime. One notable consequence of this phenomenon is the appearance of cosmological singularities during the Big Bang and the existence of black holes.
A significant challenge for GR did appear when it was faced with the problem of explaining the late-time cosmic accelerated expansion. Evidence from observations of type Ia supernovae [15, 16, 17], large-scale structure observations, and measurements of the cosmic microwave background (CMB) anisotropies from the Wilkinson Microwave Anisotropy Probe (WMAP) [18], and of the Planck satellite [19] highlighted a limitation in GR’s ability to fully describe and comprehend the dynamics of the Universe at cosmic scales during its later stages. This failure of GR prompted the exploration of alternative theories of gravity and the consideration of additional factors or components in the gravitational field equations, such as dark energy or the addition of a cosmological constant ($\Lambda$) in action ($S_{EH}$), to reconcile observations with theoretical predictions [4]. Hence, a more general gravitational framework is required to explain the gravitational dynamics across various scales, ranging from the Solar System to galaxies and the large scale Universe.
In pursuit of a more comprehensive understanding of gravity that aligns with the observational evidence, a plethora of modified gravity theories have been proposed, such as Scalar-tensor theories [20, 21, 22, 23, 24, 25], Tensor-Vector-Scalar (TeVeS) [26], Dvali-Gabadadze-Porrati (DGP) gravity [27], Einstein-Gauss-Bonnet gravity [28], brane-world gravity [29], Einstein-Aether theory [30], Eddington-inspired Born-Infeld (EiBI) gravity [31, 32] etc.
A class of gravitational theories, known as $f(R)$ gravity, arises through a straightforward extension of the Einsetin-Hilbert action $S_{EH}$ by replacing $R$ with an arbitrary functions of the Ricci scalar $R$ [33]. The geometrical structures of the $f(R)$ gravity were able to explain the accelerated cosmic expansion [34], and also the flat rotation curves of galaxies, without introducing dark matter [35]. Even though it failed when subjected to Solar-System tests [36, 37, 38], $f(R)$ gravity could still be a valuable approach to the foundational framework for a “parameterized post-Friedmann” description of linear phenomena and could draw parallels with the parameterized post-Newtonian framework for small-scale tests of gravity.
An alternative method for extending the Einstein-Hilbert action involves postulating the presence of a non-minimal coupling between geometry and matter, and it leads to the $f(R,L_{m})$ gravity [39]. For the various astrophysical and cosmological implications of this theory see [40, 41, 42, 43, 44, 45, 46, 47, 48, 49]. Another similar approach is based on the inclusion of a non-minimal coupling between geometry, described by the Ricci scalar $R$, and the trace of the energy-momentum tensor $T$, giving rise to $f(R,T)$ gravity [50]. A more comprehensive exploration of this theory is available in the detailed investigations presented in [51, 52, 53, 54, 55, 56, 57, 58, 59, 60]. In all these extended theories, the gravitational dynamics is described by more general functions of the curvature scalar, matter Lagrangian, and the trace of momentum-energy tensor, respectively, which allows for obtaining a broader range of gravitational behaviours going beyond the predictions of GR. For a detailed review of modified gravity and its implications see [61, 62, 63, 64, 33, 65, 66, 67, 68, 69, 70, 71].
GR is based solely on the metric and on the Riemannian curvature tensor to define gravity. However, within the broader context of metric-affine geometry, gravity is not limited to curvature alone; it can also be mediated by two additional geometric quantities, torsion and non-metricity, respectively.
In the context of the Riemannian geometry, the torsion tensor faces a severe limitation. Specifically, due to the symmetry of the Christoffel symbols, the torsion tensor is restricted to zero, that is, $T^{\mu}_{\rho\lambda}=0$. In an interesting extension of Riemann geometry, in the Weitzenböck space [72], the torsion tensor is non-zero ($T^{\mu}_{\rho\lambda}\neq 0$), and the Riemann curvature tensor is zero, leading to a spacetime characterized by flat geometry, endowed with a significant property known as absolute parallelism, or teleparallelism. The applications of Weitzenböck-type spacetimes in physics were pioneered by Einstein to introduce a unified teleparallel theory, unifying electromagnetism and gravity [73]. In the teleparallel approach, the fundamental characteristic is the replacement of the metric $g_{\mu\nu}$, which serves as the primary physical variable that describes gravitational properties, with a set of tetrad vectors $e^{i}_{\mu}$. Torsion, originating from the tetrad fields, can be employed as a comprehensive descriptor of the gravitational effects, thus replacing curvature with torsion. This leads to the theory known as the teleparallel equivalent of general relativity (TEGR) [74, 75, 76], which was extended to the $f(T)$ gravity theory.
In teleparallel or $f(T)$ type gravity theories, torsion exactly compensates curvature, resulting in a flat spacetime. A notable advantage of $f(T)$ gravity theory lies in its second-order field equations, which differentiates it from the $f(R)$ gravity, which, within the metric approach, is described by fourth-order field equations. [77]. The applications of $f(T)$ gravity theories have been extensively explored in the study of astrophysical processes, and in cosmology. Significantly, these theories are extensively used to provide an alternative explanation for large-scale structure, the late-time accelerating expansion of the Universe, thus eliminating the need to introduce dark energy [78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99].
The third geometric formulation of gravitational theories is based on the non-metricity $Q$ of the metric [100]. Geometrically, this quantity captures the variation in the length of a vector during parallel transport. Moreover, it offers the advantage of covariantizing conventional coordinate calculations in general relativity. In the framework of symmetric teleparallel gravity, the associated energy-momentum density is fundamentally the Einstein pseudotensor, transformed into a true tensor. In the context of gravitational actions containing non-metricity, the action $S_{STEGR}=(-1/2k)\int\sqrt{-g}Qd^{4}x+S_{m}$, which substitutes the curvature scalar with the non-metricity, is at the basic of a theory called the Symmetric Teleparallel Equivalent of General Relativity (STEGR) [101]. The extension of the symmetric teleparallel gravity led to the formulation of the $f(Q)$ gravity theory, also known as Coincident General Relativity [102] or nonmetric gravity. In this theory, the connection is flat and torsionless. These conditions lead to a connection that is purely inertial, differing from the Levi-Civita connection through a general linear gauge transformation. Furthermore, the torsionless condition simplifies the connection to $Y^{\alpha}_{\;\;\mu\beta}=(\partial x^{\alpha}/\partial\xi^{\lambda})\partial_{\mu}\partial_{\beta}\xi^{\lambda}$ for some arbitrary $\xi^{\lambda}$. This crucial outcome indicates that the connection can be entirely removed through a diffeomorphism. Consequently, the $\xi^{\lambda}$ fields emerge as Stückelberg fields, restoring this gauge symmetry [102].
In exploring extensions of symmetric teleparallel gravity, recent studies have considered the characteristics of gravitational wave propagation. An analysis particularly of the speed and polarization of gravitational waves [103] has remarkably extended the results obtained in general relativity, unveiling consistent speeds and polarizations.
In another line of research, in [104], a derivation of the exact propagator for the most general infinite-derivative, even-parity, and generally covariant theory within symmetric teleparallel spacetimes was presented. This approach involves decomposing the action, containing the non-metricity tensor and its contractions, into terms involving the metric and a gauge vector field.
Further insights emerged from the study of new general relativistic type solutions in symmetric teleparallel gravity theories [105]. The investigation of the gravitational wave propagation in a Minkowski spacetime revealed that all gravitational waves propagate with the speed of light. The Noether symmetry approach played a key role in classifying possible quadratic, first-order derivative terms of the non-metricity tensor in the framework of symmetric teleparallel geometry [106]. The cosmology of the $f(Q)$ theory and its observational constraints were considered in [107] and [108], where it was shown that the accelerating expansion is an intrinsic property of the Universe’s geometry, thus eliminating the need for exotic dark energy or additional fields. and also used a dynamical system approach. For more work, check the Refs. [109, 110, 111, 112]
Investigation of cosmological perturbations in $f(Q)$ gravity [113] revealed intriguing findings, such as the re-scaling of the Newton constant in tensor perturbations and the absence of vector contributions without vector sources being present. Notably, the scalar sector introduced two additional propagating modes, suggesting that $f(Q)$ theories add at least two extra degrees of freedom. Moreover, extending non-metric gravity by incorporating the trace of the matter-energy-momentum tensor $T$ into a general function $f(Q,T)$ has been investigated in [114, 115]. These $f(Q,T)$ gravity models have been observationally constrained as noted in [116], and some models have successfully described the accelerated expansion of the Universe [117]. Additionally, a spherically symmetric stellar system in $f(Q,T)$ gravity has been shown to satisfy all the physical conditions [118]. For more works in $f(Q,T)$ gravity, see Refs. [119, 120, 121, 122, 123, 124, 125]. Over the past two decades, numerous studies have been devoted to the geometrical and physical aspects of symmetric teleparallel gravity, with a surge in interest in recent years [126, 127, 128, 129, 130, 131, 132, 133, 103, 104, 134, 135, 136, 105, 106, 107, 108, 137, 113, 138].
Riemannian geometry represents a specific case within the broader framework of metric-affine geometry, offering a restricted perspective on gravitational dynamics. However, there exist no definitive physical principles that exclusively favor Riemannian geometry as the sole representation of gravity. Instead, metric-affine geometry presents three distinct yet physically equivalent avenues for describing gravitational phenomena. These approaches attribute the gravitational effects to the presence of non-zero curvature, non-zero torsion, or non-zero non-metricity within a given geometric framework. Together, these descriptions constitute the geometric trinity of General Relativity [139, 138]. It is essential to investigate all three approaches equally to gain a comprehensive understanding of gravity.
The coupling between the gravitational field and matter fields defines the dynamics in spacetime. In GR, the minimal coupling principle dictates that matter theories formulated in flat Minkowski space are seamlessly extended to incorporate gravitational interactions by replacing the flat metric and partial derivatives with the curved metric and covariant derivatives. This principle holds as long as the matter fields are coupled solely to the metric and its determinant without involving derivatives of the metric. In teleparallel gravity, for the electromagnetic potential, the presence of torsion introduces additional terms in the Maxwell action, violating the expected equivalence with GR [140]. Similarly, fermionic fields are affected by torsion, further challenging the minimal coupling principle. However, in symmetric teleparallel gravity, the scenario shifts. The minimal coupling principle remains intact even in the presence of non-metricity [140]. For electromagnetic fields and fermions alike, non-metricity does not interfere with the standard coupling prescriptions, ensuring compatibility with GR. In essence, while the symmetric teleparallel theory maintains equivalence with GR in the presence of matter fields, teleparallel theories diverge from this equivalence, underscoring the nuanced interplay between gravity and matter within these distinct frameworks.
The flat $\Lambda$CDM model generally aligns well with observations, but recent data indicate possible discrepancies. These include variations in the measured values of the Hubble constant $H_{0}$, and the amplitude of matter fluctuations $\sigma_{8}$, when different methods are used. Additionally, some anomalies arise when comparing the model’s theoretical predictions, based on the best-fit cosmological parameters, with actual observations. These potential inconsistencies encourage the investigation of extensions of the $\Lambda$CDM model. The well-known discrepancy between $H_{0}$, measured by the SH0ES collaboration using local distance ladder measurements from Type Ia supernovae $(H_{0}=73\pm 1\,{\rm km/s/Mpc})$ [141, 142], and the value inferred by the Planck collaboration from observations of temperature and polarization anisotropies in the Cosmic Microwave Background (CMB) radiation, assuming a $\Lambda$CDM cosmology $(H_{0}=67.4\pm 0.5\,{\rm km/s/Mpc})$ [19], has reached a statistical significance exceeding $5\sigma$. Unless this discrepancy is due to systematic errors, an intriguing possibility is that the Hubble tension could indicate new physics beyond the standard $\Lambda$CDM model of cosmology.
The main objective of our study is to generalize symmetric teleparallel gravity by incorporating the matter Lagrangian in the Lagrangian density of the $f(Q)$ theory, thus obtaining the $f(Q,L_{m})$ theory. The corresponding action could describe minimal and non-minimal couplings between geometry and matter. After introducing the basic action of the model, we obtain the general system of field equations by varying the action with respect to the metric. We also investigate the conservation problem of the matter energy-momentum tensor and show that, in the present theory, it is not conserved. Furthermore, we investigate the cosmic evolution for the case of a flat Friedmann-Lemaitre-Robertson-Walker metric by first obtaining the generalized Friedmann equations. Furthermore, we consider two particular gravitational models corresponding to two distinct forms of the function $f\left(Q,L_{m}\right)$. The predictions of the two gravitational models are compared with two distinct observational datasets. Our results unravel the intricate dynamics of the Universe within this extended gravitational framework.
The paper is structured as follows. Section 2.1 provides the geometrical background for building the general theory, which is followed by the derivation of the field equation using the metric formalism (Section 2.2). Furthermore, in Section 2.4, we obtain the energy-momentum tensor balance equation. In Section 3, the cosmic evolution of the Universe governed by $f(Q,L_{m})$ gravity is investigated. Two specific expressions for the arbitrary function $f\left(Q,L_{m}\right)$ are considered in Section 4.2 and Section 4.3, respectively. The model parameters are constrained using MCMC in Section 4.4. A concise summary of the work is presented in Section 5. The detailed calculations for obtaining the Friedmann equations are provided in Appendix 5.
2 Field equations of $f(Q,L_{m})$ gravity
This section first provides a concise overview of the geometric foundations underlying gravitational theories, which are based on the existence of a general line element in spacetime. Then we introduce the action of the $f(Q,L_{m})$ gravitational theory, and, with the help of the variational principle, we obtain the corresponding gravitational field equation, offering new insights into the understanding of gravitational phenomena within this geometric framework. Furthermore, we explore the non-conservation of the matter-energy momentum tensor, highlighting the impact of the coupling between matter Lagrangian and geometry.
2.1 Geometric Preliminaries
Once the definition of a metric is provided, the geometric interpretation of gravity is given by the Riemann tensor
$R^{\alpha}_{\;\;\beta\mu\nu}=\partial_{\mu}Y^{\alpha}_{\;\;\nu\beta}-\partial_{\nu}Y^{\alpha}_{\;\;\mu\beta}+Y^{\alpha}_{\;\;\mu\lambda}Y^{\lambda}_{\;\;\nu\beta}-Y^{\alpha}_{\;\;\nu\lambda}Y^{\lambda}_{\;\;\mu\beta},$ (2.1)
and of its contractions. The Riemann tensor is constructed with the help of an affine connection. The general form of the affine connection $Y^{\alpha}_{\;\;\mu\nu}$ consists of three parts: a symmetric part known as the Levi-Civita connection $\Gamma^{\alpha}_{\;\;\mu\nu}$, a contortion tensor $K^{\alpha}_{\;\;\mu\nu}$ describing the anti-symmetric part, and the disformation tensor $L^{\alpha}_{\;\;\mu\nu}$, accounting for the presence of non-metricity,
$Y^{\alpha}_{\;\;\mu\nu}=\Gamma^{\alpha}_{\;\;\mu\nu}+K^{\alpha}_{\;\;\mu\nu}+L^{\alpha}_{\;\;\mu\nu}.$ (2.2)
The torsion-free Levi-Civita connection $\Gamma^{\alpha}_{\;\;\mu\nu}$ is equivalent to the 2nd order Christoffel symbol in terms of the metric; it preserves the inner product of the various tangent vectors when a vector is parallelly transported, and it is defined according to
$\Gamma^{\alpha}_{\;\;\mu\nu}=\frac{1}{2}g^{\alpha\lambda}(\partial_{\mu}g_{\lambda\nu}+\partial_{\nu}g_{\lambda\mu}-\partial_{\lambda}g_{\mu\nu}).$ (2.3)
The contortion tensor $K^{\alpha}_{\;\;\mu\nu}$ is represented in terms of torsion tensor $T^{\alpha}_{\;\;\mu\nu}$ as
$K^{\alpha}_{\;\;\mu\nu}=\frac{1}{2}(T^{\alpha}_{\;\;\mu\nu}+T_{\mu\;\;\nu}^{\;\;\alpha}+T_{\nu\;\;\mu}^{\;\;\alpha}).$ (2.4)
The torsion tensor characterizes the deviation of a connection from symmetry, which indicates that parallel transport around a closed loop does not necessarily bring a vector back to its original position.
The disformation tensor $L^{\alpha}_{\;\;\mu\nu}$ describes the overall expansion or contraction of spacetime. When a vector is parallelly transported, its magnitude changes along its path. The variation of the length is measured by the non-metricity tensor,
$L^{\alpha}_{\;\;\mu\nu}=\frac{1}{2}(Q^{\alpha}_{\;\;\mu\nu}-Q^{\;\;\alpha}_{\mu\;\;\nu}-Q^{\;\;\alpha}_{\nu\;\;\mu}).$ (2.5)
The non-metricity tensor $Q_{\alpha\mu\nu}$ is defined according to
$Q_{\alpha\mu\nu}=\nabla_{\alpha}g_{\mu\nu}=\partial_{\alpha}g_{\mu\nu}-Y^{\beta}_{\;\;\alpha\mu}g_{\beta\nu}-Y^{\beta}_{\;\;\alpha\nu}g_{\mu\beta}.$ (2.6)
To construct a boundary term in the action of the metric-affine gravity theories, we need a non-metricity conjugate, known as the superpotential $P^{\alpha}_{\;\;\mu\nu}$, defined as [114]
$P^{\alpha}_{\;\;\mu\nu}=-\frac{1}{2}L^{\alpha}_{\;\;\mu\nu}+\frac{1}{4}(Q^{\alpha}-\tilde{Q}^{\alpha})g_{\mu\nu}-\frac{1}{4}\delta^{\alpha}_{\;\;(\mu}Q_{\nu)}.$ (2.7)
Here, $Q^{\alpha}=Q^{\alpha\;\;\mu}_{\;\;\mu}$ and $\tilde{Q}^{\alpha}=Q_{\mu}^{\;\;\alpha\mu}$ are the non-metricity vectors. The non-metricity scalar can be obtained by contracting the superpotential tensor with the non-metricity tensor,
$Q=-Q_{\lambda\mu\nu}P^{\lambda\mu\nu}.$ (2.8)
The non-metricity scalar $Q$ describes the deviation of the manifold geometry from isotropy and can be thought of as a measure of how much the volume of a parallelly transported object changes as it moves through spacetime.
2.2 The variational principle and the field equation
The dynamics of a physical system are studied by using the action principle. The action for the $f(Q,L_{m})$ modified gravity takes the following form
$S=\int f(Q,L_{m})\sqrt{-g}d^{4}x,$ (2.9)
where $\sqrt{-g}$ is the determinant of the metric, and $f(Q,L_{m})$ is an arbitrary function of non-metricity scalar $Q$ and of the matter Lagrangian $L_{m}$.
By varying the action with respect to the metric tensor, we obtain the gravitational field equation, which describes how spacetime geometry responds to the presence of matter and energy. Hence, we first obtain
$\delta S=\int\bigr{[}(f_{Q}\delta Q+f_{L_{m}}\delta L_{m})\sqrt{-g}+f\delta\sqrt{-g}\bigl{]}d^{4}x.$ (2.10)
Here, $f_{Q}=\partial f(Q,L_{m})/\partial Q$ and $f_{L_{m}}=\partial f(Q,L_{m})/\partial L_{m}$.
The variation of $Q$ is given by [114]
$\delta Q=2P_{\alpha\nu\rho}\nabla^{\alpha}\delta g^{\nu\rho}-(P_{\mu\alpha\beta}Q_{\nu}^{\;\;\alpha\beta}-2Q^{\alpha\beta}_{\;\;\;\mu}P_{\alpha\beta\nu})\delta g^{\mu\nu}.$ (2.11)
The energy-momentum tensor $T_{\mu\nu}$ of the matter is defined as [9]
$T_{\mu\nu}=-\frac{2}{\sqrt{-g}}\frac{\delta(\sqrt{-g}L_{m})}{\delta g^{\mu\nu}}=g_{\mu\nu}L_{m}-2\frac{\partial L_{m}}{\partial g^{\mu\nu}},$ (2.12)
The variation of the determinant of the metric tensor is
$\delta\sqrt{-g}=-\frac{1}{2}\sqrt{-g}g_{\mu\nu}\delta g^{\mu\nu}.$ (2.13)
From Eqs. (2.11), (2.12), and (2.13) it follows that Eq. (2.10) can be written as
$\delta S=\int\bigr{[}\bigr{(}f_{Q}(2P_{\alpha\nu\rho}\nabla^{\alpha}\delta g^{\nu\rho}-(P_{\mu\alpha\beta}Q_{\nu}^{\;\;\alpha\beta}-2Q^{\alpha\beta}_{\;\;\;\mu}P_{\alpha\beta\nu})\delta g^{\mu\nu})$ (2.14)
$+\frac{1}{2}f_{L_{m}}(g_{\mu\nu}L_{m}-T_{\mu\nu})\delta g^{\mu\nu}\bigl{)}-\frac{1}{2}fg_{\mu\nu}\delta g^{\mu\nu}\bigl{]}\sqrt{-g}d^{4}x.$
After applying the boundary conditions and integrating the first term in Eq. (2.14) becomes $-2\nabla^{\alpha}(f_{Q}\sqrt{-g}P_{\alpha\mu\nu})\delta g^{\mu\nu}$. Equating the metric variation of the action to zero, we obtain the field equation of the $f(Q,L_{m})$ gravity
$\frac{2}{\sqrt{-g}}\nabla_{\alpha}(f_{Q}\sqrt{-g}P^{\alpha}_{\;\;\mu\nu})+f_{Q}(P_{\mu\alpha\beta}Q_{\nu}^{\;\;\alpha\beta}-2Q^{\alpha\beta}_{\;\;\;\mu}P_{\alpha\beta\nu})+\frac{1}{2}fg_{\mu\nu}=\frac{1}{2}f_{L_{m}}(g_{\mu\nu}L_{m}-T_{\mu\nu}).$ (2.15)
For $f(Q,L_{m})=f(Q)+2\,L_{m}$, it reduces to the field equation of $f(Q)$ gravity (as seen in [143])
$\frac{2}{\sqrt{-g}}\nabla_{\alpha}(\sqrt{-g}f_{Q}\,P^{\alpha}_{\;\;\mu\nu})+f_{Q}\,q_{\mu\nu}+\frac{1}{2}f(Q)g_{\mu\nu}=-T_{\mu\nu},$ (2.16)
where $q_{\mu\nu}=P_{\mu\alpha\beta}Q_{\nu}^{\;\;\alpha\beta}-2Q^{\alpha\beta}_{\;\;\;\mu}P_{\alpha\beta\nu}$. Furthermore, the field equation (2.15) can also be reduced to the STEGR.
In the mixed tensor representation, the field equation (2.15) is given by
$\frac{2}{\sqrt{-g}}\nabla_{\alpha}(f_{Q}\sqrt{-g}P^{\alpha\mu}_{\;\;\;\;\nu})+f_{Q}\,P^{\mu}_{\;\;\alpha\beta}Q_{\nu}^{\;\;\alpha\beta}+\frac{1}{2}\delta^{\mu}_{\;\;\nu}\,f=\frac{1}{2}f_{L_{m}}(\delta^{\mu}_{\;\;\nu}L_{m}-T^{\mu}_{\;\;\nu}).$ (2.17)
Using the Lagrange multiplier method with constraints $T^{\alpha}_{\;\;\beta\gamma}=0$ and $R^{\alpha}_{\;\;\beta\mu\nu}=0$, the action (2.9) reads as
$S=\int\bigr{[}f(Q,L_{m})\sqrt{-g}+\lambda_{\alpha}^{\;\;\beta\gamma}\,T^{\alpha}_{\;\;\beta\gamma}+\xi_{\alpha}^{\;\;\beta\mu\nu}\,R^{\alpha}_{\;\;\beta\mu\nu}\bigl{]}d^{4}x.$ (2.18)
The variation of the Lagrange multipliers is given as
$\delta(\lambda_{\alpha}^{\;\;\beta\gamma}\,T^{\alpha}_{\;\;\beta\gamma})=2\,\lambda_{\alpha}^{\;\;\beta\gamma}\,\delta Y^{\alpha}_{\;\;\beta\gamma},$ (2.19)
$\delta(\xi_{\alpha}^{\;\;\beta\mu\nu}\,R^{\alpha}_{\;\;\beta\mu\nu})$ $=\xi_{\alpha}^{\;\;\beta\mu\nu}\big{[}\nabla_{\mu}(\delta\,Y^{\alpha}_{\;\;\nu\beta})-\nabla_{\nu}(\delta\,Y^{\alpha}_{\;\;\mu\beta})\big{]}$ (2.20)
$=2\,\xi_{\alpha}^{\;\;\nu\beta\mu}\nabla_{\beta}(\delta Y^{\alpha}_{\;\;\mu\nu})\simeq 2(\nabla_{\beta}\xi_{\alpha}^{\;\;\nu\beta\mu})\delta Y^{\alpha}_{\;\;\mu\nu}.$ (2.21)
Varying now the action (2.18) with respect to the connection gives
$\delta\,S=\int\left(4\sqrt{-g}\,f_{Q}\,P^{\mu\nu}_{\;\;\;\;\alpha}+H_{\alpha}^{\;\;\mu\nu}+2\,\nabla_{\beta}\xi_{\alpha}^{\;\;\nu\beta\mu}+2\,\lambda_{\alpha}^{\;\;\mu\nu}\right)d^{4}x\,\delta Y^{\alpha}_{\;\;\mu\nu}.$ (2.22)
Here $H_{\alpha}^{\;\;\mu\nu}$ is the hypermomentum density defined as
$H_{\alpha}^{\;\;\mu\nu}=\sqrt{-g}f_{L_{m}}\frac{\delta L_{m}}{\delta Y^{\alpha}_{\;\;\mu\nu}}.$ (2.23)
In the action variation, we introduce two covariant derivatives $\nabla_{\mu}\nabla_{\nu}$ to eliminate the Lagrange multiplier coefficients with the anti-symmetry property of $\mu$ and $\nu$. Then, the field equation becomes
$\nabla_{\mu}\nabla_{\nu}\Bigl{(}4\sqrt{-g}\,f_{Q}\,P^{\mu\nu}_{\;\;\;\;\alpha}+H_{\alpha}^{\;\;\mu\nu}\Bigl{)}=0.$ (2.24)
2.3 The Klein-Gordon equation
When a scalar field is coupled to the Ricci scalar in a non-minimal way, within the Lagrangian framework, it significantly alters the Einstein field equations, and the Klein-Gordon equation. These modifications have important implications for the overall dynamics of cosmological evolution. In the presence of a non-minimally coupled scalar field the Klein-Gordon equation is given by [144, 145]
$\left(\Box+m_{0}^{2}+\xi R\right)\phi=0,$ (2.25)
where $\phi$ is the scalar field, $\Box=\nabla_{\mu}\nabla^{\mu}$, $R$ is the Ricci scalar, $m_{0}$ denotes the mass of the scalar field particle, and $\xi$ is a dimensionless coupling constant. For $\xi=0$, Eq. (2.25) reduces to the standard form of the Klein-Gordon equation in the presence of a minimal coupling, $\left(\Box+m_{0}^{2}\right)\phi=0$.
The $f(Q,L_{m})$-field equation (2.15) can be rewritten in a covariant form, similar to the standard Einstein gravitational field equations as (see [146] for the detailed calculation),
$f_{Q}{G}_{\mu\nu}+\frac{1}{2}g_{\mu\nu}\left(f-f_{Q}Q\right)+2f_{QQ}\left(\partial_{\alpha}Q\right)P^{\alpha}{}_{\mu\nu}=\frac{1}{2}f_{L_{m}}(g_{\mu\nu}L_{m}-T_{\mu\nu}).$ (2.26)
In the STGR limit with $f(Q)=-Q+2L_{m}$, the left hand side of Eq. (2.26) reduces to the Einstein tensor $G_{|mu\nu}$, which only depends on the metric of the spacetime manifold.
By introducing the notations,
$\Delta=\frac{f_{L_{m}}}{f_{Q}},\,\delta=\frac{f}{f_{L_{m}}},\,\Sigma=\frac{f}{f_{Q}},\Psi_{\alpha}=2\frac{f_{QQ}}{f_{Q}}\partial_{\alpha}Q,$ (2.27)
Eq. (2.26) can be reformulated as
$G_{\mu\nu}+\frac{1}{2}g_{\mu\nu}\left(\Sigma-Q\right)+\Psi_{\alpha}P^{\alpha}_{\,\,\,\mu\nu}=\frac{1}{2}\Delta\left(g_{\mu\nu}L_{m}-T_{\mu\nu}\right).$ (2.28)
By taking the trace of Eq. (2.26), and after systematic algebraic simplifications, we obtain the Ricci scalar of the $f(Q,L_{m})$ gravity as,
$R=\frac{1}{2}\Delta\left(T-4L_{m}\right)+2\left(\Sigma-Q\right)+\partial_{\alpha}Q\left(Q^{\alpha}-\tilde{Q^{\alpha}}\right)\frac{\partial}{\partial Q}\log{f_{Q}}.$ (2.29)
Substituting Eq. (2.29) into Eq. (2.25) results in the modified Klein-Gordon equation in the $f(Q,L_{m})$ gravity,
$\left(\Box+m_{eff}^{2}\right)\phi=0,$ (2.30)
where we have denoted,
$m_{eff}^{2}=m_{0}^{2}+\xi\left[\frac{1}{2}\Delta\left(T-4L_{m}\right)+2\left(\Sigma-Q\right)+\partial_{\alpha}Q\left(Q^{\alpha}-\tilde{Q^{\alpha}}\right)\frac{\partial}{\partial Q}\log{f_{Q}}\right].$ (2.31)
Thus, $m_{eff}$ represents the effective mass of the scalar field in $f(Q,L_{m})$ gravity. The scalar field interacts not only with its own mass $m_{0}$, but also with the non-metricity and matter lagrangian, as described by the additional terms in the equation. This kind of generalization of the Klein-Gordon equation allows for rich scalar field dynamics, and can give some novel insights into the explanation of phenomena such as cosmic acceleration, inflation, or dark energy.
2.4 Energy-momentum tensor balance equations
The covariant derivative of $\omega^{\mu}_{\;\;\nu}$ is given by
$\nabla_{\mu}\,\omega^{\mu}_{\;\;\nu}=D_{\mu}\,\omega^{\mu}_{\;\;\nu}-\frac{1}{2}Q_{\rho}\,\omega^{\rho}_{\;\;\nu}-L^{\lambda}_{\;\;\mu\nu}\,\omega^{\mu}_{\;\;\lambda},$ (2.32)
where $D_{\mu}$ is the covariant derivative with respect to the Levi-Civita connection. The covariant derivative of the field equation (2.17) is
$D_{\mu}\Bigr{[}\frac{1}{2}f_{L_{m}}(\delta^{\mu}_{\nu}\,L_{m}-T^{\mu}_{\;\;\nu})\Bigl{]}=\frac{1}{2}\partial_{\nu}f+D_{\mu}(f_{Q}P^{\mu}_{\;\;\alpha\beta}Q_{\nu}^{\;\;\alpha\beta})+D_{\mu}\Bigr{[}\frac{2}{\sqrt{-g}}\nabla_{\alpha}(f_{Q}\sqrt{-g}P^{\alpha\mu}_{\;\;\;\;\nu})\Bigr{]},$ (2.33)
From Eq. (2.32) $D_{\mu}$ can be expressed as $D_{\mu}=\nabla_{\mu}+\frac{1}{2}Q_{\mu}+L^{\rho}_{\;\;\mu\nu}$. Thus, Eq. (2.33) becomes
$\frac{1}{2\sqrt{-g}}\nabla_{\alpha}\nabla_{\mu}H_{\nu}^{\;\;\alpha\mu}-\frac{1}{2}f_{L_{m}}D_{\mu}\,T^{\mu}_{\;\;\nu}=\frac{1}{2}f_{Q}\,\partial_{\nu}Q\ +\nabla_{\mu}(f_{Q}P^{\mu}_{\;\;\alpha\beta}Q_{\nu}^{\;\;\alpha\beta})\ +\frac{1}{2}Q_{\mu}(f_{Q}P^{\mu}_{\;\;\alpha\beta}Q_{\nu}^{\;\;\alpha\beta})\ +\\ L^{\rho}_{\;\;\mu\nu}(f_{Q}P^{\mu}_{\;\;\alpha\beta}Q_{\rho}^{\;\;\alpha\beta})+\frac{2}{\sqrt{-g}}L^{\rho}_{\;\;\mu\nu}\nabla_{\alpha}(\sqrt{-g}f_{Q}\,P^{\alpha\mu}_{\;\;\;\;\rho})+\frac{1}{\sqrt{-g}}Q_{\mu}\nabla_{\alpha}(\sqrt{-g}f_{Q}\,P^{\alpha\mu}_{\;\;\;\;\nu}).$ (2.34)
The detailed calculations are shown in [114], and they lead to
$D_{\mu}\,T^{\mu}_{\;\;\nu}=\frac{1}{f_{L_{m}}\sqrt{-g}}\left[\nabla_{\alpha}\nabla_{\mu}H_{\nu}^{\;\;\alpha\mu}-2\,Q_{\mu}\,\nabla_{\alpha}(f_{Q}\sqrt{-g}P^{\alpha\mu}_{\;\;\;\;\nu})\right].$
To simplify the above equation, we introduce the tensor $A^{\mu}_{\;\;\alpha}$ and define Eq. (2.24) such that
$\nabla_{\mu}\Bigl{(}4\sqrt{-g}\,f_{Q}\,P^{\mu\nu}_{\;\;\;\;\alpha}+H_{\alpha}^{\;\;\mu\nu}\Bigl{)}=\sqrt{-g}A^{\nu}_{\;\;\alpha}.$ (2.35)
Then the covariant derivative of the RHS of Eq. (2.35) is
$\nabla_{\nu}(\sqrt{-g}A^{\nu}_{\;\;\alpha})=\sqrt{-g}\nabla_{\nu}\,A^{\nu}_{\;\;\alpha}+\frac{\sqrt{-g}}{2}Q_{\nu}\,A^{\nu}_{\;\;\alpha}=0.$ (2.36)
Eq. (2.4) simplifies by the combination of Eqs. (2.35) and (2.36) as
$D_{\mu}\,T^{\mu}_{\;\;\nu}=\frac{1}{f_{L_{m}}}\left[\frac{2}{\sqrt{-g}}\nabla_{\alpha}\nabla_{\mu}H_{\nu}^{\;\;\alpha\mu}+\nabla_{\mu}\,A^{\mu}_{\;\;\nu}-\nabla_{\mu}\left(\frac{1}{\sqrt{-g}}\nabla_{\alpha}H_{\nu}^{\;\;\alpha\mu}\right)\right]=B_{\nu}\neq 0.$ (2.37)
From Eq. (2.37), it follows that the matter energy-momentum tensor is not conserved in the $f(Q,L_{m})$ gravity theory. The non-conservation tensor $B_{\nu}$ is a function of dynamical variables like $Q$, $L_{m}$, and the thermodynamic parameters.
In a broader context, dissipative processes pose significant challenges when reconciling cosmic microwave background radiation (CMBR) and large-scale structure (LSS). In [55], the cosmological and solar system consequences of a class of models with geometry-matter coupling were investigated. The findings of this work suggest that these models often exhibit inconsistent behaviour as compared to observational data. This inconsistency may potentially manifest and amplify when extended to cosmological scales at both galactic and extragalactic levels. In particular, incompatibility with CMBR or LSS appears to be a model-dependent phenomenon. However, the study in [55] reveals that some or all of these inconsistencies can be mitigated through meticulous fine-tuning of model parameters.
At larger scales, specifically galactic and extra-galactic levels, the non-minimal matter coupling with geometry introduces intriguing implications. The observed flattening of galaxy rotation curves considered a dynamically generated effect, is attributed to the non-minimal coupling [147, 148]. The non-conservation of the energy-momentum tensor leads to a deviation from geodesic motion, explaining the observed deviation between measured rotation velocity and classical predictions. Moreover, a specific type of non-minimal matter coupling with geometry is shown to mimic the presence of dark matter in galaxy clusters. In [149], they explore this phenomenon in the context of the Abell cluster A586, demonstrating its potential extension to a larger sample of galaxy clusters. Adding to the complexity of the physical behaviour, dissipative processes play a distinctive role in the evolution of radio galaxies, as discussed in [150].
If we consider matter as a perfect fluid described by its pressure $p$ and energy density $\rho$, the energy-momentum tensor can be defined as
$T^{\mu}_{\;\;\nu}=(\rho+p)u_{\nu}\,u^{\mu}+p\,\delta^{\mu}_{\nu},$ (2.38)
where $u^{\mu}$ denotes the four-velocity of the fluid. Following [136] we have
$\dot{\rho}+3\,H(\rho+p)=B_{\mu}\,u^{\mu}.$ (2.39)
The continuity equation presented above deviates noticeably from the standard form, incorporating additional terms on the right-hand side (RHS) that account for the deviations from the geodesic motion. In this context, the source term, denoted by $B_{\mu}\,u^{\mu}$, is associated with the generation or dissipation of energy. When $B_{\mu}\,u^{\mu}=0$, the system obeys the energy conservation law of standard gravity. On the contrary, if $B_{\mu}\,u^{\mu}$ takes nonzero values, the energy transfer processes become dominant.
The momentum conservation equation, which describes the movement of massive particles [114, 136], is expressed as
$\frac{d^{2}x^{\mu}}{ds^{2}}+\Gamma^{\mu}_{\alpha\beta}u^{\alpha}u^{\beta}=\frac{h^{\mu\nu}}{\rho+p}(B_{\nu}-D_{\nu}\,p)=F^{\mu},$ (2.40)
where $h^{\mu\nu}$ represents the projection operator, defined as $h^{\mu\nu}=g^{\mu\nu}+u^{\mu}u^{\nu}$. The equation of motion exhibits a notable departure from the geodesic motion of the massive particles. An additional force, $F^{\mu}$, emerges as a consequence of the coupling between $Q$ and $L_{m}$. This coupling introduces a non-gravitational influence, leading to deviations from the trajectories determined by the standard geodesic motion of general relativity thus influencing the dynamical evolution of massive particles.
3 Cosmological evolution of FLRW universe in $f(Q,L_{m})$ gravity
In the present Section, we will investigate, in a general framework, the cosmological implications of the $f(Q,L_{m})$ gravity theory. By considering a flat Friedmann-Lemaitre-Robertson-Walker geometry, the generalized Friedmann equations are derived. The cosmological evolution equations do contain some extra-terms, coming from the presence of the nonmetricity and geometry matter coupling, which generate an effective density of pressure, which can be interpreted as representing geometric dark energy. The general form of the energy balance equation is also obtained. The de Sitter limiting behaviour of the cosmological models is also investigated.
3.1 The Friedmann equations
To study the cosmological evolution in $f(Q,L_{m})$ gravity, we assume that the Universe is described by a flat Friedmann-Lemaître-Robertson-Walker (FLRW) geometry, with the spacetime interval of the form
$ds^{2}=-dt^{2}+a^{2}(t)(dx^{2}+dy^{2}+dz^{2}),$ (3.1)
where $a(t)$ is the scale factor [^1] Here, we assume the Lapse function as $N(t)=1$.. We define the rate of expansion of the universe as $H=\frac{\dot{a}}{a}$. Let us also assume that the Universe is filled with a perfect fluid. We adopt the expressions $L_{m}=-\rho$, or $L_{m}=p$ for the Lagrangian density of the cosmic matter. Hence, in the comoving frame, the non-zero components of the energy-momentum tensor are given by $T^{\mu}_{\nu}=(-\rho,p,p,p)$.
Using the FLRW metric, the field equations Eq. (2.15) give the two generalized Friedmann equations (the detailed calculations are presented in the Appendix A)
$3H^{2}=\frac{1}{4f_{Q}}\bigr{[}f-f_{L_{m}}(\rho+L_{m})\bigl{]},$ (3.2)
$\dot{H}+3H^{2}+\frac{\dot{f_{Q}}}{f_{Q}}H=\frac{1}{4f_{Q}}\bigr{[}f+f_{L_{m}}(p-L_{m})\bigl{]}.$ (3.3)
For $f(Q,L_{m})=f(Q)+2\,L_{m}$ the Friedmann equations reduces to $f(Q)$ [151, 152], it can further be simplified to STEGR. By subtracting Eqs. (3.2) and (3.3), we obtain
$\frac{d}{dt}(f_{Q}\,H)=\frac{f_{L_{m}}}{4}(p+\rho).$ (3.4)
The generalized expression of the deceleration parameter is obtained as
$q$ $=-1-\frac{\dot{H}}{H^{2}}$
$=\frac{1}{4f_{Q}H^{2}}\left(2Qf_{Q}+4\dot{f_{Q}}H-f-f_{L_{m}}(p-L_{m})\right)-1.$ (3.5)
With the help of Eq. (3.2), Eq. (3.3) can be rewritten as
$2\dot{H}+3H^{2}=\frac{1}{4f_{Q}}\left[f+f_{L_{m}}\left(\rho+2p-L_{m}\right)\right]-2\frac{\dot{f}_{Q}}{f_{Q}}H.$ (3.6)
Thus, we can reformulate the generalized Friedmann equations of the $f\left(Q,L_{m}\right)$ gravity theory in the form
[…]
where we have introduced the effective energy density and pressure, defined as
$\rho_{eff}=\frac{1}{4f_{Q}}\bigr{[}f-f_{L_{m}}(\rho+L_{m})\bigl{]},$ (3.8)
and
$p_{eff}=2\frac{\dot{f}_{Q}}{f_{Q}}H-\frac{1}{4f_{Q}}\left[f+f_{L_{m}}\left(\rho+2p-L_{m}\right)\right],$ (3.9)
respectively. Eq. (3.7) allow to formulate the generalized effective conservation equation of the $f\left(Q,L_{m}\right)$ gravity theory as
$\dot{\rho}_{eff}+3H\left(\rho_{eff}+p_{eff}\right)=0.$ (3.10)
Using Eq. 2.27 we can represent the effective energy density and pressure as
$\rho_{eff}=\frac{1}{4}\Delta\left[\delta-\left(\rho+L_{m}\right)\right],$ (3.11)
and
$p_{eff}=2\frac{\dot{f}_{Q}}{f_{Q}}H-\frac{1}{4}\Delta\left[\delta+\left(\rho+2p-L_{m}\right)\right],$ (3.12)
respectively. Then the conservation equation (3.10) can be reformulated as
$\dot{\rho}$ $+$ $3H\left(\rho+p\right)=\frac{1}{\Delta}\frac{d}{dt}\left[\Delta\left(\delta-L_{m}\right)\right]+3H\left\{8\frac{\dot{f}_{Q}}{f_{Q}}\frac{H}{\Delta}-\left[\left(1+\frac{\dot{\Delta}}{\Delta}\right)\rho+p\right]\right\}=\Gamma.$
The function $\Gamma$ describes the non-conservation level of the present modified gravity theory. If $\Gamma>0$, the energy of the particles increases due to the energy transfer of matter to the gravitational field. The case $\Gamma<0$ can be interpreted as describing particle decay due to the matter-geometry coupling.
From Eqs. (3.7) we also obtain the expression of the deceleration parameter as
$q$ $=$ $\frac{1}{2}+\frac{3}{2}\frac{p_{eff}}{\rho_{eff}}$ (3.13)
$=$ $\frac{1}{2}+6\frac{2\dot{f}_{Q}H-(1/4)\left[f+f_{L_{m}}\left(\rho+2p-L_{m}\right)\right]}{f-f_{L_{m}}\left(\rho+L_{m}\right)}.$
The Universe enters into an accelerating phase when $q<0$, or $p_{eff}<-\rho_{eff}/3$. This gives the condition that must be satisfied by the function $f$ and its derivatives to describe an accelerated expansion
$012\dot{H}f_{Q}+f-f_{L_{m}}\left(\rho+L_{m}\right)>$ 0 . (3.14)
To compare the theoretical results with the cosmological observations, we introduce an independent variable redshift $z$ instead of the usual time variable $t$, defined as $a=\frac{1}{1+z}$, where we have used a normalization of the scale factor by imposing $a(0)=1$. Thus, we can replace the derivatives with respect to the time with the derivatives with respect to the redshift using the relation
$\frac{d}{dt}=-(1+z)H(z)\frac{d}{dz}.$ (3.15)
Moreover, the redshift dependence of the deceleration parameter is given by
$q(z)=-1+(1+z)\frac{H^{\prime}(z)}{H(z)}.$ (3.16)
3.2 The de Sitter solution
As a first step in considering explicit theoretical models, we consider the problem of the existence of a de-sitter-type vacuum solution of the cosmological field equations. The de Sitter solution corresponds to $p=0$, $\rho=0$ and $H=H_{0}={\rm constant}$, respectively. For a vacuum de Sitter type Universe, Eq. (3.4) gives $\dot{f_{Q}}=0$, and further results in $f_{Q}=F_{0}$, where $F_{0}$ is a constant.
The condition $f_{Q}=F_{0}$ is satisfied for any $Q$, when we have [114, 136]
$f(Q)=F_{0}Q+2\Lambda,$ (3.17)
where $\Lambda$ is an integration constant. In the vacuum de Sitter phase, the first field equation (3.2) reduces to the form
$3H_{0}^{2}=\frac{6F_{0}H_{0}^{2}+2\Lambda}{4F_{0}}.$ (3.18)
One can also write the above equation as
$H_{0}=\sqrt{\frac{\Lambda}{3F_{0}}}.$ (3.19)
Hence, the $f(Q,L_{m})$ theory admits the de Sitter type evolution in the limiting case of a vacuum Universe. As can be easily calculated, for the de-Sitter solution, we have $q=-1$ and $\omega=-1$, respectively.
4 Cosmological models
In this Section, we will explore various cosmological models based on the $f(Q,L_{m})$ gravity theory. The models are determined by specific choices for the functional form of $f\left(Q,L_{m}\right)$. To keep our analysis as general as possible, we will assume that the matter in the Universe obeys an equation of state given by $p=(\gamma-1)\rho$ where $p$ is the pressure and $\rho$ is the energy density, $1\leq\gamma\leq 2$. For $\gamma=4/3$, this linear relationship between pressure and energy density describes the behaviour of the radiation in the early Universe, characterized by high density, as well as, for $\gamma=1$, in the present Universe, when the matter density is low. However, we will begin our presentation by describing the data sets and the statistical methods used to compare the models with observations.
4.1 Data and methodology of MCMC analyses
The present Section presents the observational datasets that constrain the $f(Q,L_{m})$ modified gravity theory with the MCMC methodology. We perform a Bayesian statistical analysis based on MCMC tools, using the Emcee module under the Python environment, to establish bounds over the model parameters. To find the maximum of the likelihood function for each data set, we use the following priors: $H_{0}:[60,80]$, $\alpha:[0,0.5]$, $\beta:[-3000,-2000]$, and $\gamma:[1,2]$. We also perform our analysis by combining the samples $OHD+SN+BAO$.
4.1.1 Datasets
  • Hubble parameter: The values of the Hubble parameter $H(z)$ are usually derived from the so-called differential age of galaxies (DAH) methodology. As the Hubble parameter can be estimated at a redshift $z$ from $H(z)=\frac{-1}{1+z}\frac{dz}{dt}$, $dz/dt$ can be obtained from measurements of massive and very slowly evolving galaxies, dubbed Cosmic Chronometers (CC). We use 31 points compiled in [153], where we take them as uncorrelated with each other.
  • Type Ia supernovae (SN): We use measurements of the most recent Pantheon$+$ dataset [154, 155], consisting of 1701 light curves from 1550 Type Ia supernovae. This is an improvement from the previous Pantheon data set, where, in particular, there is an increase in the low redshift range. This data set includes measurements of Cepheid hosts from the SH0ES collaboration, composed of 77 points. The theoretical estimate for the distance modulus $\mu(z)=m_{b}-M_{B}$ is
    $\mu_{th}=5\log_{10}D_{L}(z)/\text{Mpc}+25,$ (4.1)
    with $D_{L}(z)$ being the luminosity distance.
  • Baryonic Acoustic Oscillations (BAO): Finally, we constrain our model using BAO [156, 157, 158]. To obtain the BAO constraints, we use the acoustic scale $l_{A}=\pi\frac{d_{A}(z_{d})}{r_{s}(z_{*})}$, where $d_{A}(z)=\int^{z}_{0}\frac{dz^{\prime}}{H(z^{\prime})}$ is the angular diameter distance in the comoving coordinates and $r_{s}$ is the sound horizon determined as $r_{s}(z_{d})=\int_{z_{d}}^{\infty}\frac{c_{s}(z^{\prime})}{H(z)}$ at the drag epoch $z_{d}$ with sound speed $c_{s}(z)$. Here, $D_{V}(z_{BAO})$ is the dilation scale $D_{V}(z)=\left[\frac{d_{A}(z)^{2}cz}{H(z)}\right]^{1/3}$. Finally, we obtain the BAO constraints $\frac{d_{A}(z_{d})}{D_{V}(z_{BAO})}$.
4.1.2 Statistical analysis
To perform the MCMC sampling, we consider the chi-squared functions $\chi^{2}_{SN}$ and $\chi^{2}_{OHD}+\chi^{2}_{SN}+\chi^{2}_{BAO}$ obtained by minimizing the corresponding log-likelihood function $\mathcal{L}=\exp(-\chi^{2}/2)$. The best-fit values and their uncertainties at 68% confidence level (CL) for the datasets are reported in Table 1.
Model Data $H_{0}$ $\alpha$ $\beta$ $\gamma$ $q_{0}$ $z_{t}$
A SN $72.454\pm 0.093$ $0.1231\pm 0.0042$ $-2484.96\pm 0.10$ $1.0313^{+0.0085}_{-0.031}$ $-0.44^{+0.040}_{-0.053}$ $0.46^{+0.11}_{-0.05}$
OHD+SN+BAO $72.527\pm 0.095$ $0.1191\pm 0.0024$ $-2484.952\pm 0.097$ $1.0046^{+0.0013}_{-0.0045}$ $-0.48^{+0.022}_{-0.025}$ $0.56^{+0.04}_{-0.03}$
B SN $72.456\pm 0.094$ $0.1228^{+0.0036}_{-0.0032}$ $-2457.873\pm 0.094$ $1.046^{+0.011}_{-0.047}$ $-0.44^{+0.03}_{-0.04}$ $0.47^{+0.08}_{-0.04}$
OHD+SN+BAO $72.546\pm 0.093$ $0.1167\pm 0.0024$ $-2457.858\pm 0.099$ $1.0188^{+0.046}_{-0.018}$ $-0.49^{+0.03}_{-0.02}$ $0.56^{+0.05}_{-0.07}$
Table 1: Confidence-level constraints on the investigated models; we have considered $SN$ and $OHD+SN+BAO$.
It is useful to estimate how preferred the proposed models are in comparison with the standard $\Lambda$CDM one. We then incorporate statistical criteria, the Akaike Information Criterion (AIC) and Bayesian Information Criterion (BIC), defined as [159]
$\mathrm{AIC}=\chi^{2}_{\mathrm{min}}+2d,\quad\mathrm{BIC}=\chi^{2}_{\mathrm{min}}+d\ln N,$ (4.2)
where $d$ is the number of free parameters, and $N$ is the total size of the data.
In this criterion, if the difference in AIC value between a given model and the best one ($\Delta$AIC) is less than 4, both models are equally supported by the data. For $\Delta$AIC values in the interval $104<\Delta AIC<$ 10, the data still support the given model but less than the preferred one. For $\Delta AIC>10$, the observations do not support the given model.
Similarly, BIC discriminates between models as follows: For $\Delta BIC<2$, there is no appreciable evidence against the model. If $62<\Delta BIC<$ 6, there is modest evidence against the considered model. For the interval $106<\Delta BIC<$ 10, the evidence against the candidate model is strong, and even stronger evidence against it exists in the data when $\Delta BIC>10$. For details, check Table 2.
4.2 Model A: $f=-\alpha\,Q+2L_{m}+\beta$ with $L_{m}=p$
As a first example of a cosmological model, we consider the functional form of $f$ as represented by $f=-\alpha\,Q+2L_{m}+\beta$ with $L_{m}=p$, where $\alpha$ and $\beta$ are constants. Hence, for this particular $f(Q,L_{m})$ model with $L_{m}=p=(\gamma-1)\rho$, the Friedmann equations reduces to
$3H^{2}=-\frac{\beta}{2\alpha}+\frac{\rho}{\alpha},$ (4.3)
$2\,\dot{H}+3H^{2}=-3H^{2}(\gamma-1)-\frac{\beta\gamma}{2\alpha}.$ (4.4)
The system of Eqs. (4.3) and (4.4) does admit a de Sitter type solution for a vacuum Universe, with $H=H_{0}={\rm constant}$, corresponding to $3H_{0}^{2}=-\beta/2\alpha$, which requires that either $\alpha$ or $\beta$ are negative. The effective densities and pressures are given by
$\rho_{eff}=\frac{\rho}{\alpha}-\frac{\beta}{2\alpha},\ p_{eff}=3H^{2}(\gamma-1)+\frac{\beta\gamma}{2\alpha},$ (4.5)
leading to the energy balance equation
$\dot{\rho}+3H\left(\rho+p\right)=0,$ (4.6)
Hence, in this model, the energy-momentum tensor of the matter is conserved.
Further, using the relation $\frac{1}{H}\frac{dH}{dt}=\frac{dH}{d\,lna}$, the above equations have an exact solution for $H$, which takes the form
$H(z)=\left[\frac{(6H_{0}^{2}\alpha+\beta)(1+z)^{3\gamma}-\beta}{6\alpha}\right]^{\frac{1}{2}},$ (4.7)
where $H(0)=H_{0}$ is the present Hubble parameter. It should be noted that the positivity condition on $H(z)$ in Eq. (4.7) gives us the constraints on model parameters for the priors to be set in MCMC analysis.
4.2.1 Statistical analysis of Model A
The 1D posterior distributions and 2D confidence level contours at $1\sigma$ CL and $2\sigma$ CL are presented in Model A in Fig. 1.
Figure 1: The corner plot for the parameter space $(H_{0},\alpha,\beta,\gamma)$ with their $1\sigma$ and $2\sigma$ confidence levels for model A in $f(Q,L_{m})$ gravity.
According to our results, we obtain the $\chi^{2}$ for $SN$ and $OHD+SN+BAO$ in the case of model A. We have $\Delta AIC=2.13$ $3.4$, $\Delta BIC=1.57$ and $8.87$ for $SN$ and $OHD+SN+BAO$. We find that our model A is well supported by the observational data $SN$ and, on average, supported by the combined set $OHD+SN+BAO$.
4.2.2 Cosmological evolution: Model A
The variations as a function of the redshift of the energy density, deceleration parameter, and the effective equation of state for Model A of $f(Q,L_{m})$ gravity are represented in Figs. 2, 3, and 4, respectively. The deceleration parameter $q(z)$ shows a significant dependence on the numerical values of the model parameters. In the redshift range $z\in(-1,0.6)$, the model can reproduce well the results of the standard $\Lambda$CDM model. The deceleration parameter for model A at higher redshifts takes much larger positive values than for the $\Lambda$CDM case, indicating a decelerating evolution followed by a quicker transition to the accelerating phase. However, the model we consider enters the accelerating phase with $q<0$ at the redshift $0.46$ and $0.56$ for the $SN$ and $OHD+SN+BAO$ [160, 161, 162]. The comparison of the deceleration parameter variation of Model A with the $\Lambda$CDM paradigm shows a qualitative similarity between the two models. Cosmological phases with a de Sitter-type expansion with $q=-1$ can also be obtained at late times.
In addition, the energy density is an increasing function of redshift $z$. The increase is almost linear for small redshifts, and it is almost independent of the numerical values of model parameters. However, a significant dependence on the model parameter can be seen at higher redshifts. The $f(Q,L_{m})$ Model A can provide a viable alternative explanation for matter dynamics, as the plot shows.
The plots in Fig. 4 depict the behaviour of the effective EoS parameter versus the redshift $z$ of Model A for different sets of observational data. According to the observations, $\gamma=4/3$ represents the radiation phase, and $\gamma=1$ corresponds to the matter-dominated (non-relativistic) phase. The first model here depicts the quintessence era of the universe at present, $-1<\omega<0$ and approaches towards $\Lambda$CDM, i.e. $\omega=-1$ at late times. However, the current function $f(Q,L_{m})$ cannot fully mimic the standard cosmic evolution in a quantitative manner. However, it produces identical qualitative results.
Figure 2: The behaviour of the energy density $\overline{\rho}=\rho/3H_{0}^{2}$ as a function of redshift using the best-fit values of model parameters for model A with $1\sigma$ levels.
Figure 3: The behaviour of the deceleration parameter as a function of redshift using the best-fit values of model parameters for model A with $1\sigma$ levels. The dashed black curve represents the evolution of the $q$ in the standard $\Lambda$CDM cosmological model with $\Omega_{m}=0.3089$ and $\Omega_{\Lambda}=0.6911$, obtained from the Planck data.
Figure 4: The behaviour of the effective equation of state parameter using the best-fit values of model parameters for model A with $1\sigma$ levels.
4.3 Model B: $f=-\alpha\,Q+(2L_{m})^{2}+\beta$ with $L_{m}=p$
As a second cosmological model, we consider the functional form as $f(Q,L_{m})=-\alpha\,Q+(2L_{m})^{2}+\beta$ with $L_{m}=p$, where $\alpha>0$ and $\beta$ are constants. For the specific functional form, the Friedmann equations are reduced to
$3H^{2}=-\frac{2}{\alpha}(1-\gamma^{2})\rho^{2}-\frac{\beta}{2\alpha},$ (4.8)
$2\dot{H}+3H^{2}=-\frac{\beta}{2\alpha}-\frac{(\gamma-1)(\beta+6\alpha H^{2})}{2\alpha(\gamma+1)}.$ (4.9)
Hence, for this model, the effective energy density and pressure can be defined as
$\rho_{eff}=-\frac{2}{\alpha}(1-\gamma^{2})\rho^{2}-\frac{\beta}{2\alpha},$ (4.10)
and
$p_{eff}=\frac{\beta}{2\alpha}+\frac{(\gamma-1)(\beta+6\alpha H^{2})}{2\alpha(\gamma+1)}.$ (4.11)
For the energy balance equation, we obtain
$\dot{\rho}+\frac{3\gamma}{\gamma+1}H\rho=0,$ (4.12)
which can be integrated to give the density-scale factor dependence in the form
$\rho(a)=\rho_{0}a^{-3\gamma/(\gamma+1)},$ (4.13)
where $\rho_{0}$ is an arbitrary integration constant. Eq. (4.12) can be reformulated as
$\dot{\rho}+3\gamma H\rho=\frac{3\gamma^{2}}{\gamma+1}H\rho=\Gamma,$ (4.14)
which shows that $\Gamma$, the matter non-conservation rate, is proportional to the factor $H\rho$. In the present model $\Gamma>0$, and thus it describes energy transfer from the gravitational field to matter. Still, in the limit of very low densities matter creation processes can be safely ignored.
The expression of the function $H(z)$ for this model is obtained as
$H(z)=\left[\frac{(6H_{0}^{2}\alpha+\beta)(1+z)^{\frac{6\gamma}{1+\gamma}}-\beta}{6\alpha}\right]^{\frac{1}{2}}.$ (4.15)
Figure 5: The corner plot for the parameter space $(H_{0},\alpha,\beta,\gamma)$ with their $1\sigma$ and $2\sigma$ confidence levels for Model B in $f(Q,L_{m})$ gravity.
4.3.1 Statistical analysis
The 2D confidence level contours at $1\sigma$ CL and $2\sigma$ CL for Model B are presented in Fig. 5. Similarly, we obtain the following results for Model B: $\chi^{2}=1609.95$, and $\chi^{2}=1632.15$. Hence, using the above information, we obtain $\Delta AIC=2.05$ $2.9$, $\Delta BIC=1.49$ $8.4$ for $SN$ and $OHD+SN+BAO$. Consequently, Model B is also similarly supported by the observational data.
4.3.2 Cosmological evolution: Model B
The energy density in Fig. 6 is a monotonically increasing function of the redshift for all adopted numerical values of the model parameters. The increase is almost linear for small redshifts. However, a dependence on the model parameter can be seen at higher redshifts. This particular $f(Q,L_{m})$ model predicts a lower energy density at high and low redshifts compared to the previous model.
Figure 6: The behaviour of the energy density $\overline{\rho}=\rho/3H_{0}^{2}$ as a function of redshift using the best-fit values of model parameters for model B with $1\sigma$ levels.
In this model, the deceleration parameter provides a good qualitative concordance with the predictions of the standard cosmology. Moreover, the behaviour of $q$ overlaps at $1\sigma$ level for $SN$ and $OHD+SN+BAO$ sets. Up to redshifts of around $z\sim 1$, the deceleration parameter is roughly a constant in the range $q\in(0.2,0.5)$, and the Universe is decelerating. Furthermore, the Universe began to accelerate, and after a short cosmological time interval, it entered an accelerating phase at the redshift $0.47$ and $0.56$ for $SN$ and $OHD+SN+BAO$, for the considered $f(Q,L_{m})$ gravity model. The second model in Fig. 8 depicts the quintessence era of the universe at present, $-1<\omega<0$ and approaches towards $\Lambda$CDM, i.e. $\omega=-1$ at late times. Thus, in the present mode, we get the overlapping and tighter constraints for both datasets, $SN$ and $OHD+SN+BAO$.
Figure 7: The behaviour of the deceleration parameter as a function of redshift using the best-fit values of model parameters for model B with $1\sigma$ levels. The dashed black curve represents the evolution of the $q$ in the standard $\Lambda$CDM cosmological model with $\Omega_{m}=0.3089$ and $\Omega_{\Lambda}=0.6911$, obtained from the Planck data.
      Model       Data       $\chi^{2}$       $AIC$       $\Delta\,AIC$       $BIC$       $\Delta\,BIC$
      $\Lambda$CDM       SN       $1609.9$       $1615.90$       00       $1638.21$       00
      OHD+SN+BAO       $1631.217$       $1637.217$       00       $1653.59$       00
      A       SN       $1610.03$       $1618.03$       $2.13$       $1639.78$       $1.57$
      OHD+SN+BAO       $1632.62$       $1640.62$       $3.4$       $1662.46$       $8.87$
      B       SN       $1609.95$       $1617.95$       $2.05$       $1639.70$       $1.49$
      OHD+SN+BAO       $1632.15$       $1640.15$       $2.9$       $1661.99$       $8.4$
Table 2: The corresponding $\chi^{2}$ of the models for each sample and the information criteria AIC and BIC for the examined cosmological models.
Figure 8: The behavior of the effective equation of state parameter using the best-fit values of model parameters for model B with $1\sigma$ levels.
Figure 9: The behavior of the Hubble parameter $H(z)$ as a function of the redshift for model A and B in $f(Q,L_{m})$ gravity.
4.4 Model comparisons, and correlation results
Here, Fig. 9 illustrates the redshift evolution of the Hubble parameter, providing a clear comparison of the fit for both models and the $\Lambda$CDM at both low and high redshifts. The black line depicts the $\Lambda$CDM model. The error-bar plot of the distance modulus $\mu(z)$ and $\mu(z)-\mu_{\Lambda CDM}$ for the Pantheon+ data is shown in Fig. 10, with the standard $\Lambda$CDM model represented by a black solid line. Notably, both models correspond closely with the $\Lambda$CDM at low redshifts, but they start to diverge slightly after $z\approx 1$.
Figure 10: The behavior of $\mu(z)$ and $\mu(z)-\mu_{\Lambda CDM}(z)$ as a function of the redshift for Models A (blue dashed line) and B (red dotted line) in $f(Q,L_{m})$ gravity.
4.4.1 Correlation results
We further investigate the Pearson correlations among key physical parameters within the datasets. From the results presented in Fig. 13, we infer the following conclusions:
  • Fig. 13(a) shows that the Hubble parameter $H_{0}$ exhibits a strong positive linear relationship with redshift $z$, as indicated by the correlation coefficient of $0.95$. The standard deviation $\sigma$ shows a moderate positive correlation with both $H_{0}$ (0.37) and $z$ (0.39).
  • Fig. 13(b) illustrates that $z_{BAO}$ has a strong negative relation with $\sigma$ and $D(=d_{A}/D_{V})$, as indicated by the correlation coefficient of -0.75 and -0.87. The standard deviation $\sigma$ shows a strong positive correlation with $D$ (0.93).
  • Fig. 13(c) shows that $\mu$ has strong positive correlation with $z_{CMB}$ (0.83) and moderate negative correlation with $\sigma_{\mu}$ (-0.31). The correlation coefficient of -0.011 indicates a very weak negative linear relationship between $\sigma_{\mu}$ and $z_{CMB}$.
4.5 Effective mass of scalar field particles
As we have already mentioned, one of the important implications of the $f\left(Q,L_{m}\right)$ is related to the modification of the elementary particle equations at a fundamental level, leading to a generalization of the Klein-Gordon equation that explicitly contains the nonmetricity and the matter Lagrangian. The presence of nonmetricity, as well as of the nonminimal matter-geometry coupling leads to the modification of the particle mass $m_{0}^{2}$, which can be interpreted as an effective, cosmological parameters and time dependent mass.
By using te results of section 2.3, we obtain the effective mass of a scalar field particle as given by Eq. (2.31) for Model A as
$m_{eff}^{2}=m_{0}^{2}+\xi\left[3H^{2}(4-3\gamma)-\frac{3}{2}\left(\frac{\beta\gamma}{\alpha}\right)\right]=m_{0}^{2}+m_{QL_{m}}^{2}(t).$ (4.16)
The term $m_{QL_{m}}^{2}(t)$ gives a time-dependent correction to the particle mass. For $\gamma=4/3$, the effective mass of the particle is a constant, $m_{eff}^{2}=m_{0}^{2}-2\xi\beta/\alpha$, similarly to the standard Klein-Gordon case, but with the mass still modified due to the presence of the geometry-matter coupling. However, the statistical analysis of the cosmological models suggests a value of $\gamma\approx 1$, which does not favor a constant effective scalar particle mass during the cosmological evolution. The effective mass is proportional to the square of the Hubble function $H^{2}$, and as such, it is a decreasing function of the cosmological time.
The effective mass for Model B can be obtained from Eq. (2.31)as
$m_{eff}^{2}=m_{0}^{2}+\xi\left[\left(\frac{2-\gamma}{1+\gamma}\right)\left(6H^{2}+\frac{\beta}{\alpha}\right)-\frac{2\beta}{\alpha}\right]=m_{0}^{2}+m_{QL_{m}}^{2}(t).$ (4.17)
Similarly to Model A, the effective mass of the scalar particle is proportional to the square of the Hubble function and becomes a constant for $\gamma=2$, having the same value as for model A, $m_{eff}^{2}=m_{0}^{2}-2\xi\beta/\alpha$.
Since as indicated by the statistical analysis and comparison with the observational cosmological data, $\beta/\alpha<0$ and $\gamma\approx 1$ for both considered models, it follows that the sign of the correction term $m_{QL_{m}}^{2}(t)$ is solely determined by the coupling coefficient $\xi$ between the scalar field and geometry. The variation of the ratio $m_{eff}^{2}/m_{0}^{2}$ with respect to $\xi$ is represented at the present time, corresponding to $H=H_{0}$, for both Models A and B in Fig. 11.
If we impose the condition of the positivity of the effective mass, $m_{eff}^{2}\geq 0$, we obtain an important constraint on the numerical value of the coupling parameter $\xi$, namely $\xi\geq-0.273$. Physically acceptable values of $\xi$ can be thus both positive and negative in the given range. On the other hand, a negative effective mass may appear in some condensed matter systems [163] called metamaterials. Generally, an object with a negative effective mass will have an acceleration opposite to the direction of the applied force. But in the following, we will discard this type of behavior as unphysical in a cosmological context.
Figure 11: The present time ratio ($H=H_{0}$) of the effective mass squared $m_{eff}^{2}$ to the initial mass squared $m_{0}^{2}$ for different values of the parameter $\xi$. The left panel corresponds to Model A and the right panel to Model B. The black solid line corresponds to the $\xi=0$ case, while the color gradient (blue to red) represents different values of $\xi$, ranging from -1 to 1. The data used in both models are from the combined OHD+SN+BAO dataset, with the parameter values given in Table 1.
The variation of the effective mass of the scalar particles is represented as a function of the redshift in Fig. 12. The variation of $m_{eff}^{2}$ significantly depends on the values of $\xi$. For $\xi\in(-0.273,0)$, the ratio of the square of the effective mass and of the particle rest mass decreases to around 0.9. On the other hand, for $\xi>0$, the effective mass squared increases with the redshift, reaching a value of around 1.4 at a redshift of $z=3$. It is interesting to note that for both Models A and B, the variation of the effective mass is very similar at both quantitative and qualitative levels due to the common dependence on $H^{2}$ of both masses. However, higher differences of the effective mass with respect to the standard rest mass are expected at higher redshifts, and this increase of the mass due to the geometry-matter coupling effects may have some significant implications on the dynamical behavior of particles in the very early stages of the cosmological evolution.
Figure 12: The redshift variation of the ratio $m_{eff}^{2}/m_{0}^{2}$ for different values of the parameter $\xi$. The left panel corresponds to Model A, while the right panel corresponds to Model B. The black solid line corresponds to the $\xi=0$ case, while the color gradient (blue to red) represents different values of $\xi$, ranging from -1 to 1. The data used in both models are from the combined OHD+SN+BAO dataset, with the parameter values given in Table 1. We fixed the value of $m_{0}=6.583\times 10^{-22}$ GeV to obtain the corresponding plots.
5 Conclusions
In this paper, we have investigated the theoretical aspects of the third geometric description of gravity, known as the symmetric teleparallel gravity, or the $f(Q)$ gravity. From a geometric and mathematical perspective, $f(Q)$ gravity uses the Weylian extension of Riemann geometry, where the fundamental metricity condition no longer holds. The violation of the metricity condition thus becomes the source of gravitational phenomena, with the non-metricity scalar $Q$ playing a similar role to that of the Ricci scalar in general relativity.
In the present study we have introduced a novel class of theories, representing an extension of the $f(Q)$ gravity, where the non-metricity $Q$ is coupled non-minimally with the matter Lagrangian $L_{m}$. Mathematically, our analysis was conducted within the framework of the metric-affine formalism. Our theory is constructed similarly to the $f(Q,T)$ theory, but with the trace of the matter energy-momentum tensor replaced by the matter Lagrangian. Similarly to the energy-momentum tensor trace-curvature couplings, in $f(Q,L_{m})$ theory, the coupling between $Q$ and $L_{m}$ leads to the non-conservation of the matter energy-momentum tensor.
By applying a variational principle, we have derived the gravitational field equations for the $f(Q,L_{m})$ gravity theory. For particular choices of $f(Q,L_{m})$, it reduces to both $f(Q)$ and STEGR. This theory provides the freedom to explore different sets of coupling between $Q$ and $L_{m}$, and thus, the theory sheds light on the coupling mechanisms between the third, non-metric geometric description of gravity and matter, representing new avenues for further theoretical exploration. Consequently, the fundamental equations describing the cosmological evolution in $f(Q,L_{m})$ gravity are expressed in terms of an effective energy density and pressure of a purely geometric origin. But they also depend on the ordinary matter-energy and pressure components of the energy-momentum tensor, as well as on the functions $f(Q,L_{m})$, $f_{Q}(Q,L_{m})$, and $f_{L_{m}}(Q,L_{m})$.
Additionally, we have obtained the general relationship describing the non-conservation of the matter-energy-momentum tensor. The equation of motion of the particles reveals a notable departure from the geodesic motion for massive particles, specific to standard general relativity. An additional force emerges as a consequence of the coupling between $Q$ and $L_{m}$. This coupling introduces a non-gravitational effect, leading to deviations from the paths followed in the standard geodesic motion and influencing the dynamical evolution of massive particles. The investigations presented may also contribute to a better understanding of the geometrical formulation of gravity theories, particularly regarding the aspects related to the geometry-matter coupling.
The standard tests of the gravitational field theories, including general relativity, are usually performed in vacuum. These standard tests involved the deflection of light by massive objects, the perihelion precession of the planets, geodesic motion, and the Shapiro delay. In the case of the present $f\left(Q,L_{m}\right)$ theory, in the vacuum limit $L_{m}\rightarrow 0$, $f\left(Q,L_{m}\right)\equiv f(Q,0)\equiv f(Q)$. Hence, the present theory reduces in the vacuum case to the $f(Q)$ theory [101], and thus all the vacuum test of the two theories are the same. In particular, black hole solutions, and the propagation of the gravitational waves coincide in the two theories. However, important differences are expected in the study of compact objects, like, for example, neutron stars. The structure of neutron stars in the $f(Q)$ was considered in [164, 165]. In [165] it was shown that hybrid stars in the $f(Q)$ theory, satisfying a radial equation of state of the form $p_{r}=\alpha\rho-\beta$, where $\alpha$, $\beta$ are constants, can successfully model the observational characteristics of the Her X-1 star. Thus, $f(Q)$ gravity represents an attractive alternative in the description of the compact objects.
Similar investigations could be also performed in the framework of the $f\left(Q,L_{m}\right)$ gravity theory. The presence of the geometry-matter coupling leads naturally to an increase of the maximum allowable mass of compact stellar objects, and thus this coupling leads to a natural explanation of the high stellar masses of some neutron stars, which cannot be fully understood by using standard general relativity and the nuclear equations of state. The study of the astrophysical objects could thus prove to be a testing ground of the present modified gravity theory, in which the observed masses of the neutron stars may lead to strong observational constraints on the parameters of the theory, and on the functional form of $f\left(Q,L_{m}\right)$.
Another possibility of testing the gravitational theory proposed in the present work is via the study of the geodesic deviation equation, which describes how objects moving under the influence of gravitational fields recede or approach one another. From an astrophysical point of view one the geodesic deviation equation has important applications is in the study of the tidal forces, which have significant effects in the eccentric inspiralling neutron star binaries, on the star formation in galaxies, due to the increase of the gas accretion rates as a result of the tidal perturbations induced by close stellar companions, and on the evolution of superradiant scalar-field states around spinning black holes [121].
As we have already seen from the analysis of the cosmological aspects of the analysis of the $f\left(Q,L_{m}\right)$ gravity theory, the curvature-matter coupling significantly modifies the nature of the gravitational interaction. A similar modification is also expected in relation to the tidal forces, as well as in the equation of motion in the Newtonian limit of the theory. Thus, the detailed comparison of the theoretical predictions of the $f\left(Q,L_{m}\right)$ gravity theory related to the modifications of the tidal forces due to the presence of the geometry-matter coupling with the observational evidences obtained from the study of a large class of astrophysical phenomena could give some significant insights into the basic properties of the gravitational interaction, its geometric description, and to constrain the effects of the nonmetricity in the Universe.
Hence, in order to obtain a consistent gravitational theory one must consider its possibility of describing a large number of cosmological and astrophysical phenomena. Restricting the analysis of a given theory to only the cosmological (or astrophysical) framework may not provide enough evidence for its viability. Only testing the theory in various astrophysical/cosmological settings, which could be described in a consistent and non-contradictory way, with the same values of the coupling constants and of the functional form of the Lagrangian density of the theory, may give a full understanding of the theoretical and observational potential of a given theory.
An interesting effect of the matter-geometry coupling does also appear when one considers the standard evolution equations of the elementary particles. In the present study we have considered in detail the effects of the $f\left(Q,L_{m}\right)$ gravity on the Klein-Gordon equation, describing the evolution of scalar particles, in the presence of the gravitational field whose effects are described by the Ricci scalar. By expressing the Ricci scalar with the help of the field equations, we have obtained a generalization of the Klein-Gordon equation that also explicitly includes, beyond the effects of the nonmetricity, the effects of the geometry-matter coupling, described by the matter Lagrangian, the trace of the matter energy-momentum tensor, as well as the derivatives of the Lagrangian density $f\left(Q,L_{m}\right)$ with respect to $L_{m}$. All these extra effects can be combined in a single term that gives an effective contribution to the particle rest mass $m_{0}$. Hence, the modified gravity effects generate an effective mass, who may have important implications on the scalar particle evolution in the early Universe. For both considered cosmological models the effective mass is proportional to $H^{2}$, and, for the obtained values of the optimal model parameters, the sign of the effective mass is determined by the coupling parameter $\xi$ between the geometry and the scalar field. The condition of the positivity of the effective mass allows us to obtain some constraints on the value of $\xi$. The redshift variation of $m_{eff}^{2}$ is also dependent on the sign and numerical values of $\xi$, and thus the effective mass can either increase, or decrease during the cosmological evolution. This variation of the effective mass resulting from the coupling between matter and geometry, as well as from the presence of nonmetricity, could have potentially important implications for the behavior of the scalar fields in both early and late Universe, in phenomena like inflation, reheating, Big Bang Nucleosynthesis, or the recent accelerated expansion.
For the description of the dynamics of the Universe, we have adopted the homogeneous and isotropic Friedmann-Lemaitre-Robertson-Walker type metric, describing the cosmological evolution in a flat geometry. In this study, we have examined two specific classes of cosmological models by adopting some simple functional forms of $f(Q,L_{m})$. For the first case, we have assumed the simple additive Lagrangian, $f(Q,L_{m})=-\alpha Q+2L_{m}+\beta$. For this model, we have obtained a wide range of cosmological scenarios and evolution corresponding to the specific numerical values of the model parameters. These scenarios may include cosmological evolution describing both the decelerating and the accelerating expansion phases of the Universe and de Sitter-type dynamics at late times. The model $f(Q,L_{m})=-\alpha Q+2L_{m}+\beta$ can provide an effective description of cosmological data up to redshifts of around $z\approx 1$. Specifically, in this model, the Universe undergoes a rapid transition from a decelerating phase, characterized by a positive value of $q$, to an accelerating state where $q<0$. This transition can result, in its final stages, in a de Sitter-type expansion. The second model with $f(Q,L_{m})=-\alpha\,Q+(2L_{m})^{2}+\beta$ also evolves from a decelerating to an accelerating state. The nature of the cosmological evolution is heavily influenced by the numerical values of the model parameters and the specific functional form of $f$. Our fundamental finding for the specific models and for the range of the cosmological parameters we have examined indicates that the Universe initially underwent a decelerating phase in its recent evolution, followed by an accelerating phase with $q<0$, in which the Universe entered at $z=z_{crit}$, and which continued for $0<z<z_{crit}$. In the future, the Universe enter into an accelerating de Sitter-type phase, with $q=-1$. From our analysis, it follows that $H_{0}$ in the second model has higher values as compared to the first model, while the values of other model parameters decrease under the same scenario. Additionally, we find consistent DE EoS behavior with the assumption of quintessence dynamics within $1\sigma$.
(a) OHD
(b) BAO
(c) Pantheon+
Figure 13: Correlation matrix heatmap for (a) Observed Hubble Data (OHD), (b) Baryonic Acoustic Oscillations (BAO) and (c) SN (Pantheon+SHOES)
We have also compared the theoretical predictions of the $f(Q,L_{m})$ theory with the corresponding results in the standard $\Lambda$CDM cosmology. Both our considered $f(Q,L_{m})$ models align well with $\Lambda$CDM at lower redshifts. However, at higher redshifts, significant differences emerge in the behavior of the $\mu(z)$ function and the deceleration parameter, as compared to the $\Lambda$CDM model. Henceforth, in the presence of matter, the models give an acceptable description of the observational data, as well as of the $\Lambda$CDM model, but without reproducing it exactly at the present time.
Another potential application of the $f(Q,L_{m})$ theory would be to consider inflation in the presence of scalar fields, which might offer a completely new perspective on the geometrical, gravitational, and cosmological processes that significantly influenced the early dynamics of the Universe. Consequently, the predictions of the present model could lead to major differences compared to those of standard general relativity or its extensions that ignore the role of matter. These differences could impact several current areas of interest, such as cosmology, gravitational collapse, and the generation of gravitational waves. To conclude, in the present investigation, we have introduced a new version of the symmetric teleparallel theory, and we have demonstrated its theoretical consistency. This approach also motivates and encourages the exploration of further extensions within the $f(Q,L_{m})$ family of theories.
Appendix A Derivation of the Friedmann equations
The metric tensor components are given by $g_{\mu\nu}={\rm diag}\left(-1,a^{2},a^{2},a^{2}\right)$, $g^{\mu\nu}={\rm diag}\left(-1,a^{-2},a^{-2},a^{-2}\right)$, and its determinant $\sqrt{-g}=a^{3}$. For the nonmetricity tensor, have the following non-zero terms,
$Q_{011}=Q_{022}=Q_{033}=2a\dot{a},$ (A1)
[…]
$P^{011}=P^{022}=P^{033}=-\frac{\dot{a}}{a^{3}},$ (A7)
$P_{011}=P_{022}=P_{033}=a\dot{a},$ (A8)
[…]
$P_{110}=P_{101}=P_{220}=P_{202}=P_{330}=P_{303}=-\frac{a\dot{a}}{4},$ (A10)
$P^{110}=P^{101}=P^{220}=P^{202}=P^{330}=P^{303}=-\frac{\dot{a}}{4a^{3}}.$ (A11)
The non-metricity scalar $Q$ is calculated using Eq.(2.8) as
$Q=-(Q_{011}P^{011}+Q_{022}P^{022}+Q_{033}P^{033}).$ (A12)
We obtain thus $Q=6H^{2}$, where $H=\dot{a}/a$.
The energy-momentum tensor $T_{\mu\nu}$ for a perfect fluid has the components
$T_{\mu\nu}={\rm diag}(\rho,pa^{2},pa^{2},pa^{2}).$ (A13)
Evaluating the field equation (2.15) for the tt-component
[…]
gives the first generalized Friedmann equation
$3H^{2}=\frac{1}{4f_{Q}}\bigr{[}f-f_{L_{m}}(\rho+L_{m})\bigl{]}.$ (A16)
By evaluating the field equation (2.15) for the xx-component
[…]
$\frac{2}{a^{3}}\frac{\partial}{\partial t}(f_{Q}a^{3}(-a\dot{a}))-2f_{Q}\bigr{(}\frac{2\dot{a}}{a}\bigl{)}(a\dot{a})+\frac{a^{2}}{2}f=\frac{a^{2}}{2}f_{L_{m}}(L_{m}-p),$ (A18)
leads to the second generalized Friedmann equation
$\dot{H}+3H^{2}+\frac{\dot{f_{Q}}}{f_{Q}}H=\frac{1}{4f_{Q}}\bigr{[}f+f_{L_{m}}(p-L_{m})\bigl{]}.$ (A19)
Data Availability
Data sharing is not applicable to this article as no datasets were generated or analyzed during the current study.
Acknowledgments
AH expresses gratitude to Mr. Sayantan Ghosh (BITS-Pilani, Hyderabad) for insightful discussions on modified gravity, and extends thanks to Dr. Subhadip Sau (Jhargram Raj College) and Dr. Massimo Rossi (INAF/OAS Bologna) for valuable conversations regarding the Klein-Gordon equation. PKS acknowledges Science and Engineering Research Board, Department of Science and Technology, Government of India for financial support to carry out Research project No.: CRG/2022/001847 and IUCAA, Pune, India, for providing support through the visiting Associateship program. We would like to thank to the anonymous reviewer and to the journal Editor for comments and suggestions that helped us to significantly improve our work.